OpenStudy (anonymous):

Write the first trigonometric function in terms of the second for θ in the given quadrant. cos θ, sin θ; θ in quadrant II

OpenStudy (anonymous):

Do you know the definitions of sin in relation to cos and tan in relation to cot and then sec in relation to csc and cot in relation to tan?

OpenStudy (anonymous):

I know that the second set are inverses. I am doing practice problems for pre calc summer school but i have no idea what i am doing

OpenStudy (anonymous):

hey i appreciate that! thanks!! how do i express this in terms of the second quadrant?

OpenStudy (anonymous):

Now all you have to is find the equation from sin, since they said they want you to write cos in terms of sin

OpenStudy (anonymous):

Well since it's in the 2nd quadrant and the x points will be cos(theta) and y points will be sin(theta) (from the unit circle the signs of the points in the QII will be 

OpenStudy (anonymous):

-cos, +sin

OpenStudy (anonymous):

oh sweet!! so for the equation it should be I/csc(theta), but because i want the cos negative it should be csc(theta) right?

OpenStudy (anonymous):

remember they want us to use sin not csc

OpenStudy (anonymous):

They want us to write cos(x) in terms of sin(x)

OpenStudy (anonymous):

oh shoot thats right

OpenStudy (anonymous):

When x is in the 2nd quadrant

OpenStudy (anonymous):

Our equation shows us that \[\Huge \cos(\theta)= \sin(\frac{\pi}{2}-\theta)\]

OpenStudy (anonymous):

so i just have to manuver that equation into terms of sin

OpenStudy (anonymous):

Now \[\Huge \frac{\pi}{2}=90 \] degrees

OpenStudy (anonymous):

Now we have to see if we put a negative in front of sin or not

OpenStudy (anonymous):

i see how you get that

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok, well our question asked for the equation of cos in terms of sin. We got that Now it says that the point of the angle is in the second quadrant, so we might have to play around with the equation like change the pi/2 or put a negative. What we're going to do is test a point and see if this equation holds true Now remember the quadrant we're worrying about is the second one so it's all the angles between 90-180 so let's pick a point inbetween those two We'll use 120 just for random sake. try plugging in cos(120) (make sure your calculator is in deg mode) and tell me what you get

OpenStudy (anonymous):

to the first equation you put up right?

OpenStudy (anonymous):

Yeah, just put in 120 for \[\Large \theta \]

OpenStudy (anonymous):

You can do it to the equation \[\Huge \cos(\theta)= \sin(\frac{\pi}{2}-\theta)\] and see if both sides come out the same, remember that the pi/2 changes to 90

OpenStudy (anonymous):

they are the same! -.5

OpenStudy (anonymous):

Ok, so we're done (: Good job

OpenStudy (anonymous):

thanks for the help!! i really appreciate it!

OpenStudy (anonymous):

Would you mind doing me a favor? If you click on my name or my picture you can go to see detailed profile and then can you leave a testimonial just saying how I helped you out as a teacher, I'd really appreciate it